# Emergent spin polarization from $\rho$ meson condensation in rotating hadronic matter

### v1

@@ I Introduction
The study of strongly interacting matter under extreme conditions of temperature, density, and rotation is a central pursuit in high-energy nuclear physics. Non-central heavy-ion collisions in laboratories such as the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) create a novel state of matter called Quark-Gluon Plasma (QGP), endowed with immense angular momentum, leading to the strongest vortical fields known in nature [1]. Simultaneously, the dense cores of neutron stars provide a natural astrophysical setting where high baryon density and rapid rotation combine, influencing the equation of state and observable phenomena like pulsar glitches and gravitational waves [2, 3]. These diverse systems offer complementary avenues to explore the rich phase structure of quantum chormodynamics.
Among the various probes of this exotic matter, vector mesons, particularly the $\rho$ meson ($J^{PC}=1^{--}$) serve as important probes. Due to their short lifetime and strong coupling to the dense medium, their in-medium spectral properties are a sensitive indicator of chiral symmetry restoration and collective effects [4, 5]. While the condensation of pseudoscalar pions ($\pi$, $J^{PC}=0^{-+}$) has been extensively studied [6, 7, 10, 8, 9, 10], the potential Bose-Einstein Condensation (BEC) of their vector counterparts, the $\rho$ mesons, remains a significantly less explored frontier. This possibility arises in high-baryon-density environments where the $\rho$ meson chemical potential $\mu_{\rho}$ is driven to match its effective mass $m_{\rho}^{*}$, a condition which can be attained in low-energy heavy-ion collisions or inside the neutron star [12, 11]. In such regimes, a strong magnetic field can stabilize the $\rho$ condensate against the decays of $\rho$ mesons [13]. The $\rho$ condensate can couple to the background magnetic field, forming a spin-aligned, superconducting state as proposed for QCD in strong magnetic fields in Ref. [13, 14].
@@ II Formalism
where the energy of each particle in the presence of rotation is given by, $\varepsilon_{l,i}$ = $\sqrt{k_{r}^{2}+k_{z}^{2}+m_{i}^{2}}-(l+s)\omega$. The $r$ is the radial distance from the axis of the rotation, whereas $J_{\nu}$ are the Bessel functions of the first kind. The radial and longitudinal momenta are given by $k_{r}$ and $k_{z}$, respectively. The causality condition requires imposing a boundary condition at $r=R$ such that $R\omega\leq 1$ ($c=1$ in natural units). We have considered $R$ = 5 fm throughout this study. This boundary condition led to the quantization of the angular momentum as $k_{r}=\xi_{l,i}/R$ [28], with $\xi_{l,i}$ being the ithith zero of the Bessel function such that, $J_{l}(\xi_{l,i})=0$. The discretization effect of momenta affects especially the lower momenta region, and therefore, the $k_{r}$ integration faces a lower bound = $\xi_{l,i}/R=\xi_{l,i}\omega$.
@@
where $n_{q_{\uparrow}}$ and $n_{q_{\downarrow}}$ are the quark densities of spin up and spin down states, respectively. Therefore, by analogy, we can study the polarization at the hadronic level by studying their spin alignment. A deviation of $\rho_{00}$ from the statistical limit of $1/3$ indicates the presence of spin alignment induced by vorticity or other medium effects, while a nonzero $P$ measures the vector polarization of the gas. We examine the density of each spin state of the rho meson gas in both the condensate and the excited states. The spin alignment or the polarization of the particle can therefore be described by estimating the net spin density normalized by the total spin density as
where the $n_{\uparrow}$ and $n_{\downarrow}$ are the densities of up and down spin states of $\rho$ mesons, respectively. The polarization defined in Eq. (14) quantifies the net spin alignment of the $\rho$ meson ensemble along the global rotation axis and represents the vector polarization, reflecting the imbalance between the populations of the spin-up and spin-down substrates. Our formalism, based on relativistic quantum statistics with rotation, provides a robust and well-established first insight into the expected phenomena, setting the stage for future, more complex effective field theory studies.
@@ III Results and Discussion
In this work, we have considered a total density of rho meson gas as $n_{\rm tot}$ = 0.15 fm-3. This accounts for all the spin states of the $\rho$ meson, namely the spin-up, spin-down, and spin-zero components within a cylindrical rotating system of fixed radius $R$ = 5 fm. In the absence of rotation ($\omega=0$), the system is symmetric, and all three spin states possess equal densities, each contributing $0.05~\text{fm}^{-3}$. At finite rotation, this degeneracy is lifted, leading to a redistribution of density among the spin states. With an increase in rotation, the up-spin density increases, while the spin-down states decrease, keeping spin-zero states almost unchanged, such that the total density is fixed at 0.15 fm-3. Let us first look at the effect of rotation on the BEC of the $\rho$ meson gas. Eq. (II) implies that below $T_{c}$, as $T\to 0$, the condensate fraction approaches unity. This means the total density considered here ($n_{\rm tot}$ = 0.15 fm-3) is in the condensate state. With an increase in $T$, the condensate fraction decreases, and at $T=T_{c}$, the condensate vanishes, leaving behind all the particles to be in the excited state. In the left panel of Fig. 1, we show the condensate fraction as a function of temperature for three different values of $\omega$, including the non-rotating case, for which $\omega$ = 0. As expected, for all cases, the condensate fraction is zero at $T_{c}$ and increases as one goes towards low $T$, approaching unity as $T$ tends to zero. With an increase in the magnitude of rotation, the BEC transition temperature, $T_{c}$, is observed to increase, which can be understood by studying the interplay of $T$, $\omega$, and $\mu$ with the constraint of fixed number density. The right panel of Fig. 1 shows the temperature dependence of the chemical potential of the system. At zero rotation, the system reduces to the simple case of a non-rotating relativistic Bose gas for which the BEC happens when $\mu=m$. This behavior is confirmed in the right panel of Fig. 1. At zero rotation, $\mu=m=0.775$ GeV for $T<T_{c}$, and at high temperature (for $T>T_{c}$), the requirement of fixed number density results in a decrease in $\mu$, consistent with the expected thermodynamic behavior [16, 10].
@@
For a specific value of rotation, the sum of the spin density of each state at any temperature equals the total density of the $\rho$ gas considered. However, the density for each state varies, and so does the net spin density. Therefore, we estimate the polarization separately for the excited ($P_{\rm ex}$) and condensate state ($P_{\rm cond}$) to study their behaviour under BEC. Since at $\omega$ = 0 GeV, all spin states have the same density, we investigate the polarization by considering three finite values of $\omega$; $\omega$ = 0.003, 0.005, and 0.02 GeV. In the upper panel of Fig. 3, the polarization results for excited and condensate particles are shown as a function of temperature. We observe a finite polarization for the excited density, which is maximum at low temperatures and decreases with increasing temperature. The polarization increases in magnitude with a higher rotation magnitude, being larger for the case of $\omega$ = 0.02 GeV within the considered temperature range. As $T\to 0$, polarization for the excited particles is maximum, although the fraction of excited particles is also very less, as evident from Fig. 1. At very high $\omega$, the density of excited particles is negligible; moreover, the density of down states is even much smaller, and whatever the excited density is, it is mostly due to spin up states. Therefore, at high $\omega$, almost fully polarized rho gas is observed at very low $T$. On the other hand, just at the onset of BEC, the condensate has a finite polarization which increases in magnitude with an increase in the value of $\omega$. As one approaches $T\rightarrow 0$, the condensate polarization increases slowly (with increasing condensate density as in Fig. 1).
We must keep in mind that $n_{\rm ex}$ and $n_{\rm cond}$ represent different populations with potentially different polarization states. While $n_{ex}$ will get polarized due to $\omega$ and $T$ effects, $n_{\rm cond}$ can have coherent polarization just from the macroscopic occupation of the ground state. The condensate itself is described by a macroscopic wavefunction. We define its polarization $P_{\text{cond}}$ as the spin alignment of this coherent ground state.
@@
In the lower panel of Fig. 3, the polarization of the excited and condensate phases is obtained as a function of total condensate fraction. Our results show that the polarizations of both the condensate $n_{\text{cond}}$ and the excited states $n_{\text{ex}}$ increase gradually with the condensate fraction $n_{\text{cond}}/n_{\text{total}}$. This indicates that the emergence of Bose-Einstein condensation actively amplifies the global spin alignment beyond a simple population of polarized states. The condensate, once formed, acts as a source of a coherent mean field that breaks the spin-rotation symmetry. This field not only defines the condensate’s own polarization but also exerts an aligning torque on the thermal cloud, resulting in the observed concurrent increase in $P_{\text{ex}}$. The condensate fraction $n_{\text{c}}/n_{\text{tot}}$ thus serves as a direct control parameter for the overall polarization strength of the system, unifying the behavior of both condensed and non-condensed components.
@@ IV Conclusion
In this work, we present the first systematic study of the interplay between rotation, Bose-Einstein condensation, and spin polarization for $\rho$ (vector) mesons in a dense, hadronic system. Our results demonstrate that global rotation not only enhances $\rho$-meson condensation by lowering the critical chemical potential $\mu_{c}$ but also induces a strong spin alignment of the condensate. A key finding is that the onset of BEC actively amplifies this polarization, as the macroscopic wavefunction of the condensate acts as a coherent mean field that breaks spin-rotation symmetry and aligns both the condensed and thermal components. Our results have direct implications for two frontiers of physics. In the context of heavy-ion collisions, the predicted large polarization anisotropy should be observable as a characteristic modulation in the angular distribution of dilepton decays ($\rho^{0}\to e^{+}e^{-}/\mu^{+}\mu^{-}$), providing a clean, penetrating probe of the high-baryon-density phase. For neutron stars, the presence of a spin-polarized $\rho$-meson condensate would not only soften the equation of state but could also generate an anisotropic pressure, potentially leading to observable macroscopic effects such as stellar deformation.

### v2

@@ I Introduction
The study of strongly interacting matter under extreme conditions of temperature, density, and rotation is a central pursuit in high-energy nuclear physics. Non-central heavy-ion collisions in laboratories such as the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) create a novel state of matter called Quark-Gluon Plasma (QGP), endowed with immense angular momentum, leading to the strongest vortical fields known in nature [1]. Simultaneously, the dense cores of neutron stars provide a natural astrophysical setting where high baryon density and rapid rotation combine, influencing the equation of state and observable phenomena like pulsar glitches and gravitational waves [2, 3]. These diverse systems offer complementary avenues to explore the rich phase structure of quantum chromodynamics.
Among the various probes of this exotic matter, vector mesons, particularly the $\rho$ meson ($J^{PC}=1^{--}$) serve as important probes. Due to their short lifetime and strong coupling to the dense medium, their in-medium spectral properties are a sensitive indicator of chiral symmetry restoration and collective effects [4, 5]. While the condensation of pseudoscalar pions ($\pi$, $J^{PC}=0^{-+}$) has been extensively studied [6, 7, 10, 8, 9], the potential Bose-Einstein Condensation (BEC) of their vector counterparts, the $\rho$ mesons, remains a significantly less explored frontier. This possibility arises in high-baryon-density environments where the $\rho$ meson chemical potential $\mu_{\rho}$ is driven to match its effective mass $m_{\rho}^{*}$, a condition which can be attained in low-energy heavy-ion collisions or inside the neutron star [12, 11]. In such regimes, a strong magnetic field can stabilize the $\rho$ condensate against the decays of $\rho$ mesons [13]. The $\rho$ condensate can couple to the background magnetic field, forming a spin-aligned, superconducting state as proposed for QCD in strong magnetic fields in Ref. [13, 14].
@@ II Formalism
where the energy of each particle in the presence of rotation is given by, $\varepsilon_{l,i}$ = $\sqrt{k_{r}^{2}+k_{z}^{2}+m_{i}^{2}}-(l+s)\omega$. The $r$ is the radial distance from the axis of the rotation, whereas $J_{\nu}$ are the Bessel functions of the first kind. The radial and longitudinal momenta are given by $k_{r}$ and $k_{z}$, respectively. The causality condition requires imposing a boundary condition at $r=R$ such that $R\omega\leq 1$ ($c=1$ in natural units). We have considered $R$ = 5 fm throughout this study. This boundary condition led to the quantization of the angular momentum as $k_{r}=\xi_{l,i}/R$ [28], with $\xi_{l,i}$ being the ithith zero of the Bessel function such that, $J_{l}(\xi_{l,i})=0$. The discretization effect of momenta affects especially the lower momenta region, and therefore, the $k_{r}$ integration faces a lower bound = $\xi_{l,i}/R=\xi_{l,i}\omega$. It is important to emphasize that the formalism has a radial dependence through the Bessel function, $J_{\nu}(k_{r}r)$. However, to extract and interpret the genuine rotational effects and to avoid any ambiguity because of spatial variations across the radius, we remove this radial dependence by evaluating all thermodynamic quantities at a fixed value of $r$ as previously done in [28, 29, 30]. This removes any additional complications arising from spatial inhomogeneity, which can be explored in detail elsewhere. Given that the centrifugation effects will be maximum at the edge of the system, all the results here are obtained by fixing $r=R$ [30].
@@
where $n_{q_{\uparrow}}$ and $n_{q_{\downarrow}}$ are the quark densities of spin up and spin down states, respectively. Therefore, by analogy, we can study the polarization at the hadronic level by studying their spin alignment. A deviation of $\rho_{00}$ from the statistical limit of $1/3$ indicates the presence of spin alignment induced by vorticity or other medium effects, while a nonzero $P$ measures the vector polarization of the gas. Our formalism, based on relativistic quantum statistics with rotation, provides a robust and well-established first insight into the expected phenomena, setting the stage for future, more complex effective field theory studies.
@@ III Results and Discussion
In this work, we have considered a total density of rho meson gas as $n_{\rm tot}$ = 0.15 fm-3. This accounts for all the spin states of the $\rho$ meson, namely the spin-up, spin-down, and spin-zero components within a cylindrical rotating system of fixed radius $R$ = 5 fm. This value of $R$ is motivated by various experimental and theoretical estimates of characteristic system sizes, which are approximately of the order 4 to 9 fm [36, 37, 38]. In the absence of rotation ($\omega=0$), the system is symmetric, and all three spin states possess equal densities, each contributing $0.05~\text{fm}^{-3}$. At finite rotation, this degeneracy is lifted, leading to a redistribution of density among the spin states. With an increase in rotation, the up-spin density increases, while the spin-down states decrease, keeping spin-zero states almost unchanged, such that the total density is fixed at 0.15 fm-3. Let us first look at the effect of rotation on the BEC of the $\rho$ meson gas. Eq. (II) implies that below $T_{c}$, as $T\to 0$, the condensate fraction approaches unity. This means the total density considered here ($n_{\rm tot}$ = 0.15 fm-3) is in the condensate state. With an increase in $T$, the condensate fraction decreases, and at $T=T_{c}$, the condensate vanishes, leaving behind all the particles to be in the excited state. In the left panel of Fig. 1, we show the condensate fraction as a function of temperature for three different values of $\omega$, including the non-rotating case, for which $\omega$ = 0. As expected, for all cases, the condensate fraction is zero at $T_{c}$ and increases as one goes towards low $T$, approaching unity as $T$ tends to zero. With an increase in the magnitude of rotation, the BEC transition temperature, $T_{c}$, is observed to increase, which can be understood by studying the interplay of $T$, $\omega$, and $\mu$ with the constraint of fixed number density. The right panel of Fig. 1 shows the temperature dependence of the chemical potential of the system. At zero rotation, the system reduces to the simple case of a non-rotating relativistic Bose gas for which the BEC happens when $\mu=m$. This behavior is confirmed in the right panel of Fig. 1. At zero rotation, $\mu=m=0.775$ GeV for $T<T_{c}$, and at high temperature (for $T>T_{c}$), the requirement of fixed number density results in a decrease in $\mu$, consistent with the expected thermodynamic behavior [16, 10].
@@
We examine the density of each spin state of the rho meson gas in both the condensate and the excited states. The spin alignment or the polarization of the particle can therefore be described by estimating the net spin density normalized by the total spin density as
where the $n_{\uparrow}$ and $n_{\downarrow}$ are the densities of up and down spin states of $\rho$ mesons, respectively. The polarization defined in Eq. (14) quantifies the net spin alignment of the $\rho$ meson ensemble along the global rotation axis and represents the vector polarization, reflecting the imbalance between the populations of the spin-up and spin-down substrates. For a specific value of rotation, the sum of the spin density of each state at any temperature equals the total density of the $\rho$ gas considered. With rotation, there is an imbalance in the spin-up ($n_{\uparrow}$) and spin-down density ($n_{\downarrow}$), and these are separately fixed due to the constraint of fixed total density. However, these fixed populations are redistributed between the condensate and excited fractions as the temperature varies across the BEC transition. Consequently, the temperature dependence of the polarization arises solely from the evolving condensate and thermal fractions, while the total polarization remains determined exclusively by the rotation. Therefore, using Eq. (14), we estimate the polarization separately for the excited ($P_{\rm ex}$) and condensate state ($P_{\rm cond}$) to study their behaviour under BEC. Since at $\omega$ = 0 GeV, all spin states have the same density, we investigate the polarization by considering three finite values of $\omega$: $\omega$ = 0.003, 0.005, and 0.02 GeV. Although experimental estimations from global hyperon polarization [1] show that $\omega$ varies in the range from $\sim$ 0.003 to 0.015 GeV, to investigate the effect of strong rotation on the dynamics of the polarization of the condensed and excited state, we also consider $\omega$ as high as 0.02 GeV. In the upper panel of Fig. 3, the polarization results for excited and condensate particles are shown as a function of temperature. We observe a finite polarization for the excited density, which is maximum at low temperatures and decreases with increasing temperature. The polarization increases in magnitude with a higher rotation magnitude, being larger for the case of $\omega$ = 0.02 GeV within the considered temperature range. As $T\to 0$, polarization for the excited particles is maximum, although the fraction of excited particles is also very less, as evident from Fig. 1. At very high $\omega$, the density of excited particles is negligible; moreover, the density of down states is even much smaller, and whatever the excited density is, it is mostly due to spin up states. Therefore, at high $\omega$, almost fully polarized rho gas is observed at very low $T$. On the other hand, just at the onset of BEC, the condensate has a finite polarization which increases in magnitude with an increase in the value of $\omega$. As one approaches $T\rightarrow 0$, the condensate polarization increases slowly (with increasing condensate density as in Fig. 1).
@@
We must keep in mind that $n_{\rm ex}$ and $n_{\rm cond}$ represent different populations with potentially different polarization states. While $n_{ex}$ will get polarized due to $\omega$ and $T$ effects, $n_{\rm cond}$ can have coherent polarization just from the macroscopic occupation of the ground state. The condensate itself is described by a macroscopic wavefunction. We define its polarization $P_{\text{cond}}$ as the spin alignment of this coherent ground state. In the lower panel of Fig. 3, the polarization of the excited and condensate phases is obtained as a function of total condensate fraction. Our results show that the polarizations of both the condensate $n_{\text{cond}}$ and the excited states $n_{\text{ex}}$ increase gradually with the condensate fraction $n_{\text{cond}}/n_{\text{total}}$. This indicates that the emergence of Bose-Einstein condensation actively amplifies the global spin alignment beyond a simple population of polarized states. The condensate, once formed, acts as a source of a coherent mean field that breaks the spin-rotation symmetry. This field not only defines the condensate’s own polarization but also exerts an aligning torque on the thermal cloud, resulting in the observed concurrent increase in $P_{\text{ex}}$. The condensate fraction $n_{\text{c}}/n_{\text{tot}}$ thus serves as a direct control parameter for the overall polarization strength of the system, unifying the behavior of both condensed and non-condensed components. It is noteworthy to mention that we have assumed a non-interacting $\rho$ gas for our study. The interaction among the particles may lead to a modification of the chemical potential, which can shift the value of $T_{c}$ from that of a non-interacting case. Also, in a highly dense interacting medium, there is a possibility of the reduction of the in-medium mass of the vector mesons, which may lead to a decrease in $\mu_{c}$ required for the BEC transition [11, 40]. However, the mechanism due to rotation, such as the rotational enhancement of $T_{\rm c}$ and its effect on polarization, remains relatively the same. In this study, we present a foundational study using an ideal gas model to isolate and establish the novel interplay between rotation, BEC, and spin polarization. A detailed study involving an interacting medium could be done to understand the medium effect explicitly.
@@ IV Conclusion
In this work, we present the first systematic study of the interplay between rotation, Bose-Einstein condensation, and spin polarization for $\rho$ mesons in a dense, hadronic system. Our results demonstrate that global rotation not only enhances $\rho$-meson condensation by lowering the critical chemical potential $\mu_{c}$ but also induces a strong spin alignment of the condensate. A key finding is that the onset of BEC actively amplifies this polarization, as the macroscopic wavefunction of the condensate acts as a coherent mean field that breaks spin-rotation symmetry and aligns both the condensed and thermal components. Our results have direct implications for two frontiers of physics. In the context of heavy-ion collisions, the predicted large polarization anisotropy should be observable as a characteristic modulation in the angular distribution of decay daughters. When the system has high vorticity of the order of $\omega\sim$ 0.015 GeV [1], our analysis shows that the critical chemical potential required for the onset of $\rho$ meson condensation is lowered to roughly two-fifths of the $\rho$ meson mass. Such a combined condition of moderately large chemical potential and vorticity is naturally achieved in heavy-ion collisions at low centre-of-mass energies. Using the established parameterizations relating chemical potential and center-of-mass energies [41, 42], one finds that heavy-ion collisions in the range $\sqrt{s_{\rm NN}}\sim$ 3 to 10 GeV, accessible at RHIC, FAIR, and NICA can result in high chemical potential and vorticity, and may provide suitable conditions to study the effect of BEC phenomena in the $\rho$ meson polarization. For neutron stars, the presence of a spin-polarized $\rho$-meson condensate would not only soften the equation of state but could also generate an anisotropic pressure, potentially leading to observable macroscopic effects such as stellar deformation.